When 6 identical cell of no internal resistance are connected to the secondary circuit of the potentiometer,the balancing length is l,the balancing length is l/3 when some cells are wrongly connected , the number of cells connected wrongly are
 
 

Grade:12

1 Answers

Khimraj
5 years ago
According to formula
E1/E2 = l1/l2
E1 = 6
E2 = 6 - 2m
l1 = l
l2 =l/3
then
6/(6 – 2m) = 3
6 = 18 – 6m
6m = 12
so m = 2
Hope it clears.